# Better bounds on Grothendieck constants of finite orders

### v2

@@ III.3 Exact results in asymmetric scenarios
In DIB+23 this problem is reformulated into a Quadratic Unconstrained Binary Optimisation (QUBO) instance, which is then given to the solver QuBowl from RKS22 . The instance solved there involves $m_{1}=m_{2}=97$ and is bigger than the ones from DBV17 (of maximal size $m_{1}=m_{2}=92$). However, it is worth noticing that the solution of the exact instance solved in DIB+23 stands out in the landscape of all binary variables. More precisely, the exact solution is likely to be unique and its value is far above the other feasible points, so that the solver could efficiently exclude vast portions of the search space. This empirical observation is supported by two facts: instances from later stages of convergence of the Frank-Wolfe method cannot be solved by QuBowl (even within ten times as much time), and the symmetric instance $A$ of size $63\times 63$ obtained from the $E_{7}$ root system (see Table 1) cannot be solved by QuBowl either, although it is way smaller.
@@
[…]
@@
Table 2: Better lower bounds on $K_{G}\left({d}{}\right)$ compared to the state of the art (SoA) from DIB+23 ; DBV17 . They are obtained by choosing two line structures in dimension $d$ defining a scalar product matrix $P\in\mathrm{SDP}_{d}^{m_{1},m_{2}}$, and then by running our Frank-Wolfe method to obtain a good separating hyperplane $M$ with respect to the correlation polytope $\mathrm{SDP}_{1}^{m_{1},m_{2}}$ (sometimes even a facet), see Section III.1. Eventually, we compute the first level of the Lasserre hierarchy to either numerically confirm the optimality of $P$ in Eq. 3 or build a slightly better matrix $P^{\prime}\in\mathrm{SDP}_{d}^{m_{1},m_{2}}$. In Table 2b, we could solve $\mathrm{SDP}_{1}\left({M}\right)$ and $K_{G}\left({d}{}\right)\geqslant\mathrm{SDP}_{d}\left({M}\right)/\mathrm{SDP}_{1}\left({M}\right)$ immediately follows from Eq. 5. In Table 2a, we could not solve $\mathrm{SDP}_{1}\left({M}\right)$ so that the bounds shown are heuristic; see Table 1 for analytical values. Note that the correct value of the SoA for $K_{G}\left({4}{}\right)$ in DBV17 is taken from Section III B therein and not from the erroneous abstract. By monotonicity, our bound on $K_{G}\left({5}{}\right)$ propagates to all $K_{G}\left({d}{}\right)$ for $d\geqslant 5$; it therefore beats the best bounds on Grothendieck constants of finite orders up to $d=9$, where $K_{G}\left({9}{}\right)\geqslant 1.48608$ comes from BBT11 . Moreover, it also beats the value $K_{G}\left({7}{}\right)\geqslant 1.48719$ that we had found with the $E_{7}$ root system in Table 1.
@@ IV.2 Numerical lower bounds on $K_{G}\left({d}\mapsto{2}\right)$
$\mathrm{SDP}_{d}\left({M}\right)=\max\left\{\sum_{x=1}^{m_{1}}\sum_{y=1}^{m_{2}}M_{xy}\left\langle{\begin{pmatrix}a_{x,1}\\ \vdots\\ a_{x,d}\end{pmatrix}},{\begin{pmatrix}b_{y,1}\\ \vdots\\ b_{y,d}\end{pmatrix}}\right\rangle\leavevmode\nobreak\ \Bigg{|}\leavevmode\nobreak\ \forall x\in[m_{1}],\leavevmode\nobreak\ \sum_{i=1}^{d}a_{x,i}^{2}=1,\leavevmode\nobreak\ \forall y\in[m_{2}],\leavevmode\nobreak\ \sum_{i=1}^{d}b_{y,i}^{2}=1\right\}.$ (13)
@@
1.0560 1.0560 Cuboctahedron x  $A_{3}$  0.9209
6 1.0983 1.0983 Icosahedron x 0.8252
@@
4 $\frac{9}{8}$ $=1.125\hphantom{0}$  300 1.1741 120-cell $H_{4}$
5 $\frac{512}{45\pi^{2}}$ $\approx 1.1528$ 20 1.1540 1.1469 x $D_{5}$ 1.2654
@@ IV.3 Exact upper bound on $K_{G}\left({3}\mapsto{2}\right)$
$\alpha\,\eta_{1}\eta_{2}\,\mathrm{SDP}_{d}\left({M}\right)\leqslant\max\left\{\sum_{x=1}^{m_{1}}\sum_{y=1}^{m_{2}}M_{xy}\left\langle{\tilde{a}_{x}},{\tilde{b}_{y}}\right\rangle\leavevmode\nobreak\ \Big{|}\leavevmode\nobreak\ \forall x\in[m_{1}],\leavevmode\nobreak\ \tilde{a}_{x}\in B^{n},\leavevmode\nobreak\ \forall y\in[m_{2}],\leavevmode\nobreak\ \tilde{b}_{y}\in B^{n}\right\}\leqslant\mathrm{SDP}_{n}\left({M}\right),$ (18)
@@
To obtain our bound on $K_{G}\left({3}\mapsto{2}\right)$, we start from the Gram matrix $P$ obtained from a centrosymmetric polyhedron with 912 vertices, that is, 406 lines. Note that this polyhedron is different from the one used in DIB+23 and features a slightly higher shrinking factor through the following construction: starting from an icosahedron, apply four times the transformation adding the normalised centers of each (triangular) facet. Taking $v_{0}=0.8962$ and running our Frank-Wolfe algorithm, we obtain a decomposition using 7886 extreme points of $\mathrm{SDP}_{2}^{406}$ and recovering $v_{0}P$ up to a Euclidean distance of $\varepsilon\approx 2.7\times 10^{-4}$. This algorithm is strictly identical to the one in DIB+23 except for the LMO, which corresponds to the case decribed in Section III.1 when $n=2$. Following (DIB+23, , Appendix D) we can make this decomposition purely analytical by noting that the unit ball for the Euclidean norm is contained in $\mathrm{SDP}_{1}$ and therefore in $\mathrm{SDP}_{2}$. Note that, contrary to DIB+23 where the extremal points of $\mathrm{SDP}_{1}$ were automatically analytical as they only have ±$\pm 1$ elements, an extra step is needed here to make numerical extremal points of $\mathrm{SDP}_{2}$ analytical. Here we follow a rationalisation procedure very similar to the one described in (DIB+23, , Appendix A) and we obtain a value $\alpha=v_{0}/(1+\varepsilon)$ that can be used in Eq. 14 to derive the bound
@@
IV.4 Implications in Bell nonlocality
@@
Works anticipating the connection with $K_{G}\left({d}\mapsto{2}\right)$ can be traced back to 2008 PV08 ; VP08 , when Vértesi and Pál studied the realisation of Bell inequalities with real quantum mechanics and the advantage of $d$-dimensional quantum systems over real ones. This continued with the work of Briët et al. BBT11 but gained traction when the powerful result by Renou et al. RTW+21 showed that complex numbers are necessary for the complex formalism.
@@ V Discussion
Beyond the perhaps anecdotal improvement brought to their respective constants, developing such tools could unlock access to higher-dimensional structures with a size way larger than anything that numerical methods could ever consider solving. This could in turn give rise to lower bounds on Grothendieck constants of finite order strong enough to compete with the best known lower bound on the Grothendieck constant (of infinite order), namely, the one presented by Davie in 1984 Dav84 and independently rediscovered by Reeds in 1991 Ree91 , both works being unpublished. This bound reads
@@
We also note that our techniques naturally apply in the complex case, for which similar Grothendieck constants have also been defined. However, since the complex Grothendieck constants of finite order do not have known interpretations, we refrained from deriving bounds on them, although our code is already capable of dealing with this case.
@@ VI Code availability
The code used to obtain the results presented in this article is part of the Julia package BellPolytopes.jl introduced in DIB+23 and based on the FrankWolfe.jl package BCP21 ; BDH+24 . While the algorithm used to derive bounds on $K_{G}\left({d}{}\right)$ is directly available in the standard version of this package, the extension to generalised constants $K_{G}\left({d}\mapsto{n}\right)$ can be found on the branch mapsto of this repository. Installing this branch can be directly done within Julia package manager by typing add BellPolytopes#mapsto.
@@ VII Acknowledgements
The authors thank Péter Diviánszky for fruitful discussions and his help to run the branch-and-bound algorithm from DBV17 on asymmetric instances; Jim Reeds and Neil Sloane for sharing the details of the derivation of the value $45/31$ for the $E_{8}$ root system mentioned in FR94 ; Timotej Hrga, Nathan Krislock, Thi Thai Le, and Daniel Rehfeldt for trying to solve our symmetric instances with various solvers; Fernando Mário de Oliveira Filho for asking details about the proof of Eq. 14, which helped us improve its presentation; Mathieu Besançon, Daniel Brosch, Wojciech Bruzda, Dardo Goyeneche, Volker Kaibel, David de Laat, Frank Valentin, and Karol Życzkowski for inspiring discussions. This research was partially supported by the DFG Cluster of Excellence MATH+ (EXC-2046/1, project id 390685689) funded by the Deutsche Forschungsgemeinschaft (DFG). T. V. acknowledges the support of the European Union (QuantERA eDICT) and the National Research, Development and Innovation Office NKFIH (Grants No. 2019-2.1.7-ERA-NET-2020-00003 and No. K145927).

### v3

@@ III.3 Exact results in asymmetric scenarios
In DIB+23 this problem is reformulated into a Quadratic Unconstrained Binary Optimisation (QUBO) instance, which is then given to the solver QuBowl from RKS23 . The instance solved there involves $m_{1}=m_{2}=97$ and is bigger than the ones from DBV17 (of maximal size $m_{1}=m_{2}=92$). However, it is worth noticing that the solution of the exact instance solved in DIB+23 stands out in the landscape of all binary variables. More precisely, the exact solution is likely to be unique and its value is far above the other feasible points, so that the solver could efficiently exclude vast portions of the search space. This empirical observation is supported by two facts: instances from later stages of convergence of the Frank-Wolfe method cannot be solved by QuBowl (even within ten times as much time), and the symmetric instance $A$ of size $63\times 63$ obtained from the $E_{7}$ root system (see Table 1) cannot be solved by QuBowl either, although it is way smaller.
@@
  • $\mathrm{SDP}_{1}\left({M}\right)=13921227005628453160441$ (solved with QuBowl RKS23 ).
@@
Table 2: Better lower bounds on $K_{G}\left({d}{}\right)$ compared to the state of the art (SoA) from DIB+23 ; DBV17 . They are obtained by choosing two line structures in dimension $d$ defining a scalar product matrix $P\in\mathrm{SDP}_{d}^{m_{1},m_{2}}$, and then by running our Frank-Wolfe method to obtain a good separating hyperplane $M$ with respect to the correlation polytope $\mathrm{SDP}_{1}^{m_{1},m_{2}}$ (sometimes even a facet), see Section III.1. Eventually, we compute the first level of the Lasserre hierarchy to either numerically confirm the optimality of $P$ in Eq. 3 or build a slightly better matrix $P^{\prime}\in\mathrm{SDP}_{d}^{m_{1},m_{2}}$. In Table 2b, we could solve $\mathrm{SDP}_{1}\left({M}\right)$ and $K_{G}\left({d}{}\right)\geqslant\mathrm{SDP}_{d}\left({M}\right)/\mathrm{SDP}_{1}\left({M}\right)$ immediately follows from Eq. 5. In Table 2a, we could not solve $\mathrm{SDP}_{1}\left({M}\right)$ so that the bounds shown are heuristic; see Table 1 for analytical values. Note that the correct value of the SoA for $K_{G}\left({4}{}\right)$ in DBV17 is taken from Section III B therein and not from the erroneous abstract. By monotonicity, our bound on $K_{G}\left({5}{}\right)$ propagates to all $K_{G}\left({d}{}\right)$ for $d\geqslant 5$; it therefore beats the best bounds on Grothendieck constants of order up to $d=9$, where $K_{G}\left({9}{}\right)\geqslant 1.48608$ comes from BBT11 . Moreover, it also beats the value $K_{G}\left({7}{}\right)\geqslant 1.48719$ that we had found with the $E_{7}$ root system in Table 1.
@@ IV.2 Numerical lower bounds on $K_{G}\left({d}\mapsto{2}\right)$
$\mathrm{SDP}_{d}\left({M}\right)=\max\left\{\sum_{x=1}^{m_{1}}\sum_{y=1}^{m_{2}}M_{xy}\left\langle{\begin{pmatrix}a_{x,1}\\ \vdots\\ a_{x,d}\end{pmatrix}},{\begin{pmatrix}b_{y,1}\\ \vdots\\ b_{y,d}\end{pmatrix}}\right\rangle\penalty 10000\ \Bigg|\penalty 10000\ \forall x\in[m_{1}],\penalty 10000\ \sum_{i=1}^{d}a_{x,i}^{2}=1,\penalty 10000\ \forall y\in[m_{2}],\penalty 10000\ \sum_{i=1}^{d}b_{y,i}^{2}=1\right\}.$ (13)
@@
1.0560 1.0560 Cuboctahedron $\checkmark$ $A_{3}$ 0.9209
6 1.0983 1.0983 Icosahedron $\checkmark$ 0.8252
@@
4 $\frac{9}{8}$ $1.125\hphantom{0}$ 300 1.1741 120-cell $H_{4}$
5 $\frac{512}{45\pi^{2}}$ $\approx 1.1528$ 20 1.1540 1.1469 $\checkmark$ $D_{5}$ 1.2654
@@ IV.3 Exact upper bound on $K_{G}\left({3}\mapsto{2}\right)$
$\alpha\,\eta_{1}\eta_{2}\,\mathrm{SDP}_{d}\left({M}\right)\leqslant\max\left\{\sum_{x=1}^{m_{1}}\sum_{y=1}^{m_{2}}M_{xy}\left\langle{\tilde{a}_{x}},{\tilde{b}_{y}}\right\rangle\penalty 10000\ \Big|\penalty 10000\ \forall x\in[m_{1}],\penalty 10000\ \tilde{a}_{x}\in B^{n},\penalty 10000\ \forall y\in[m_{2}],\penalty 10000\ \tilde{b}_{y}\in B^{n}\right\}\leqslant\mathrm{SDP}_{n}\left({M}\right),$ (18)
@@
To obtain our bound on $K_{G}\left({3}\mapsto{2}\right)$, we start from the Gram matrix $P$ obtained from a centrosymmetric polyhedron with 912 vertices, that is, 406 lines. Note that this polyhedron is different from the one used in DIB+23 and features a slightly higher shrinking factor through the following construction: starting from an icosahedron, apply four times the transformation adding the normalised centers of each (triangular) facet. Taking $v_{0}=0.8962$ and running our Frank-Wolfe algorithm, we obtain a decomposition using 7886 extreme points of $\mathrm{SDP}_{2}^{406}$ and recovering $v_{0}P$ up to a Euclidean distance of $\varepsilon\approx 2.7\times 10^{-4}$. This algorithm is strictly identical to the one in DIB+23 except for the LMO, which corresponds to the case decribed in Section III.1 when $n=2$. Following (DIB+23, , Appendix D) we can make this decomposition purely analytical by noting that the unit ball for the Euclidean norm is contained in $\mathrm{SDP}_{1}$ and therefore in $\mathrm{SDP}_{2}$. Note that, contrary to DIB+23 where the extremal points of $\mathrm{SDP}_{1}$ were automatically analytical as they only have $\pm 1$ elements, an extra step is needed here to make numerical extremal points of $\mathrm{SDP}_{2}$ analytical. Here we follow a rationalisation procedure very similar to the one described in (DIB+23, , Appendix A) and we obtain a value $\alpha=v_{0}/(1+\varepsilon)$ that can be used in Eq. 14 to derive the bound
@@
IV.4 Implications for Bell nonlocality
@@
Works anticipating this connection with $K_{G}\left({d}\mapsto{2}\right)$ go back to 2008 PV08 ; VP08 , when Vértesi and Pál investigated the realisation of Bell inequalities in real quantum mechanics and showed that $d$-dimensional complex systems can outperform real ones. This line of research was further developed by Briët et al. BBT11 .
@@ V Discussion
Beyond the perhaps anecdotal improvement brought to their respective constants, developing such tools could unlock access to higher-dimensional structures with a size way larger than anything that numerical methods could ever consider solving. This could in turn give rise to lower bounds on finite-order Grothendieck constants strong enough to compete with the best known lower bound on the (infinite-order) Grothendieck constant, namely, the one presented by Davie in 1984 Dav84 and independently rediscovered by Reeds in 1991 Ree91 , both works being unpublished. This bound reads
@@
We also note that our techniques naturally apply in the complex case, for which similar Grothendieck constants have also been defined. However, since the complex finite-order Grothendieck constants do not have known interpretations, we refrained from deriving bounds on them, although our code is already capable of dealing with this case.
@@ VI Code availability
The code used to obtain the results presented in this article is part of the Julia package BellPolytopes.jl introduced in DIB+23 and based on the FrankWolfe.jl package BCP22 ; BDH+24 . While the algorithm used to derive bounds on $K_{G}\left({d}{}\right)$ is directly available in the standard version of this package, the extension to generalised constants $K_{G}\left({d}\mapsto{n}\right)$ can be found on the branch mapsto of this repository. Installing this branch can be directly done within Julia package manager by typing add BellPolytopes#mapsto.
@@ VII Acknowledgements
The authors thank Péter Diviánszky for fruitful discussions and his help to run the branch-and-bound algorithm from DBV17 on asymmetric instances; James A. Reeds and Neil J. A. Sloane for sharing the details of the derivation of the value $45/31$ for the $E_{8}$ root system mentioned in FR94 ; Timotej Hrga, Nathan Krislock, Thi Thai Le, and Daniel Rehfeldt for trying to solve our symmetric instances with various solvers; Fernando Mário de Oliveira Filho for asking details about the proof of Eq. 14, which helped us improve its presentation; Mathieu Besançon, Daniel Brosch, Wojciech Bruzda, Dardo Goyeneche, Volker Kaibel, David de Laat, Frank Valentin, and Karol Życzkowski for inspiring discussions. This research was partially supported by the DFG Cluster of Excellence MATH+ (EXC-2046/1, project id 390685689) funded by the Deutsche Forschungsgemeinschaft (DFG). T. V. acknowledges the support of the European Union (QuantERA eDICT, CHIST-ERA MoDIC), the National Research, Development and Innovation Office NKFIH (Grants No. 2019-2.1.7-ERA-NET-2020-00003, No. 2023-1.2.1-ERA_NET-2023-00009 and No. K145927) and the ‘Frontline’ Research Excellence Program of the NKFIH (No. KKP133827).